# DefendCLI: {Command-Line} Driven Attack Provenance Examination

### v1

@@ V-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)
TABLE VII: Attack-TTPs and Node-Level Results on Industrial Real-Time Detection
@@
TABLE VIII: Attack Scenario Checklist (Industrial)
@@ 3 Evilnum [65] Spear Phishing Python RAT
V-B GitHub Availability, Reproducibility and Real-Time Detection
Table III shows the availability (A), reproducibility (RP), and real-time detection capability (RT) of related work based on big-4 security conferences (C). The research work of DISTDET [14] and Prographer [12] are not publicly available on GitHub or other sources. Therefore, our comparative study focuses on the remaining work. The verification of availability was conducted on 7 February 2025.
V-C Licensed Commercial EDR Solutions
Table IV presents the detection capabilities of commercial EDR solutions across four key security features: anti-virus (AV), web security (WS), cyber threat intelligence (CTI), and real-time (RT) protection. For our evaluation, we obtained commercial licenses for Microsoft Defender [56], Symantec EDR [57], and Kaspersky EDR [58, 59]. Each EDR solution was updated to the latest version, ensuring the most recent threat database was used during testing.
V-D Attack Provenance Evaluation Criterion
@@
V-E Overall Evaluation
V-E1 Detection Results on DARPA E3 Benchmark
Table VI presents the threat model of the three datasets in the E3 benchmark, while Table V compares node-level detection results with state-of-the-art research. Our analysis shows that while all existing methods can detect E3 attack scenarios, their node-level results vary significantly. This discrepancy highlights the presence of substantial non-attack-related information in their alarms, leading to lower detection precision and reduced $F_{1}$ scores.
@@ V-E1 Detection Results on DARPA E3 Benchmark
V-E2 Detection Results in the Industrial Real-Time Detection
Table VIII presents the threat model for industrial real-time detection, incorporating modern APT threats relevant to 2025. These advanced threats are designed to evade modern defense strategies, enabling a more comprehensive evaluation of our solution’s detection capabilities in complex industrial environments.
@@ V-E2 Detection Results in the Industrial Real-Time Detection
V-F Ablation Evaluation
V-F1 Evaluating the Interoperability of Attack Provenance Examination Through Attack-Clause Sketch and Evidence Awareness
@@ V-F Ablation Evaluation
TABLE IX: The Use and Not Use of Community Detection and SimHash Recommender on Node-Level $F_{1}$ Score
@@
Table IX shows the node-level $F_{1}$ score with or without the use of community detection. The experimental results can support our claim that the application of community detection improves correlation accuracy for identifying interphase attacks, leading to more precise and effective threat analysis.
@@
Table IX presents the node-level $F_{1}$ score with and without the use of SimHash recommender. The results demonstrate that incorporating the SimHash recommender enhances context awareness by effectively differentiating command-line activities, leading to more precise threat analysis.
@@
V-F2 Cross-Platform Detection Reliability: Performance Across Heterogeneous OS Environments
@@
V-F3 Adaptive Detection in Dynamic Cyber Environments: Minimizing False Positives
@@
V-F4 Enhancing the Practicability of our Solution for Actionable Attack Provenance Analysis
@@
TABLE X: The Scale of Attack-Related Command-Line Detected by DefendCLI and Nodlink
@@
V-G The Improvement of Command-Line Detection Capability
Table X presents a comparison of the scale of attack-related command-line activities detected by DefendCLI and Nodlink [13]. It can be observed that our solution demonstrates superior overall effectiveness in identifying a greater number of malicious command-line activities.
$\blacksquare$ How well does DefendCLI in the detection of obfuscation compared to Nodlink? Figures 18(a) and 18(c) illustrate the detection results of both detectors in handling advanced command-line transformations and Base64-encoded shell command obfuscation.
In Figure 18(a), the string 732F5B202F67732F5B...202F67 represents a transformation of an ASCII command into its hexadecimal form as a defense evasion technique. When converted back to ASCII, it reveals the command $s/[[:blank:]]\textbackslash{}+//g$, which is used to clean up injected spaces in obfuscated shellcode designed to evade detection. In Figure 18(c), the string 2E2F676C20262E2F67...6C2026 represents a base64 obfuscated command-line for defense evasion. When decoded the content, it reveals the command $gtcache~\&\textgreater{}/dev/null~\&$, which runs a backdoor process in the background, suppressing output and ensuring persistent session activity without logging.
$\blacksquare$ How well does DefendCLI in the attack correlation with the low-frequency event compared to Nodlink? Figures 18(b), 18(d), 18(e), and 18(f) illustrate the detection results of both detectors in handling complex attack-chain correlations involving low-frequency events.
V-H Runtime Performance
Table XI presents the execution performance of DefendCLI across diverse datasets, highlighting its efficiency through parallelization. For the E3 data sets, DefendCLI runs on a 64,GB RAM, 16-core virtual machine, processing approximately 5,000 records in 6 seconds using multiprocessing. For industrial-scale, real-time detection tasks (A1, A2, A3), the system is deployed on Microsoft Azure DataLake [66], leveraging Kafka-based streaming and ETL pipelines. Under this setup, DefendCLI achieves end-to-end response times of roughly 2 seconds, including network latency, for high-volume threat detection workloads.
VI Artifact Availability
The complete artifact package accompanying this submission is available as open-source at https://anonymous.4open.science/r/DefendCLI-6CDF. This package includes the source code for our proposed method (DefendCLI), the datasets used in our evaluations, and scripts to reproduce the experiments reported in the paper. The repository is structured to facilitate artifact evaluation; please refer to the included README file for guidance.
TABLE XI: Runtime Performance

### v2

@@ IV-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)
TABLE VI: Attack-TTPs and Node-Level Results on Industrial Real-Time Detection
@@
TABLE VII: Attack Scenario Checklist (Industrial)
@@ 3 Evilnum [28] Spear Phishing Python RAT
IV-B GitHub Availability, Reproducibility and Real-Time Detection
Table II shows the availability (A), reproducibility (RP), and real-time detection capability (RT) of related work based on big-4 security conferences (C). The research work of DISTDET [5] and Prographer [36] are not publicly available on GitHub or other sources. Therefore, our comparative study focuses on the remaining work. The verification of availability was conducted on 7 February 2025.
IV-C Licensed Commercial EDR Solutions
Table III presents the detection capabilities of commercial EDR solutions across four key security features: anti-virus (AV), web security (WS), cyber threat intelligence (CTI), and real-time (RT) protection. For our evaluation, we obtained commercial licenses for Microsoft Defender [20], Symantec EDR [32], and Kaspersky EDR [14, 15]. Each EDR solution was updated to the latest version, ensuring the most recent threat database was used during testing.
IV-D Attack Provenance Evaluation Criterion
@@
IV-E Overall Evaluation
IV-E1 Detection Results on DARPA E3 Benchmark
Table V presents the threat model of the three datasets in the E3 benchmark, while Table IV compares node-level detection results with state-of-the-art research. Our analysis shows that while all existing methods can detect E3 attack scenarios, their node-level results vary significantly. This discrepancy highlights the presence of substantial non-attack-related information in their alarms, leading to lower detection precision and reduced $F_{1}$ scores.
@@ IV-E1 Detection Results on DARPA E3 Benchmark
IV-E2 Detection Results in the Industrial Real-Time Detection
Table VII presents the threat model for industrial real-time detection, incorporating modern APT threats relevant to 2025. These advanced threats are designed to evade modern defense strategies, enabling a more comprehensive evaluation of our solution’s detection capabilities in complex industrial environments.
@@ IV-E2 Detection Results in the Industrial Real-Time Detection
IV-F Ablation Evaluation
IV-F1 Evaluating the Interoperability of Attack Provenance Examination Through Attack-Clause Sketch and Evidence Awareness
@@ IV-F Ablation Evaluation
TABLE VIII: The Use and Not Use of Community Detection and SimHash Recommender on Node-Level $F_{1}$ Score
@@
Table VIII shows the node-level $F_{1}$ score with or without the use of community detection. The experimental results can support our claim that the application of community detection improves correlation accuracy for identifying interphase attacks, leading to more precise and effective threat analysis.
@@
Table VIII presents the node-level $F_{1}$ score with and without the use of SimHash recommender. The results demonstrate that incorporating the SimHash recommender enhances context awareness by effectively differentiating command-line activities, leading to more precise threat analysis.
@@
IV-F2 Cross-Platform Detection Reliability: Performance Across Heterogeneous OS Environments
@@
IV-F3 Adaptive Detection in Dynamic Cyber Environments: Minimizing False Positives
@@
IV-F4 Enhancing the Practicability of our Solution for Actionable Attack Provenance Analysis
@@
IV-G The Improvement of Command-Line Detection Capability
Table IX presents a comparison of the scale of attack-related command-line activities detected by DefendCLI and Nodlink [17]. It can be observed that our solution demonstrates superior overall effectiveness in identifying a greater number of malicious command-line activities.
$\blacksquare$ How well does DefendCLI in the detection of obfuscation compared to Nodlink? Figures 16(a) and 16(c) illustrate the detection results of both detectors in handling advanced command-line transformations and Base64-encoded shell command obfuscation.
In Figure 16(a), the string 732F5B202F67732F5B...202F67 represents a transformation of an ASCII command into its hexadecimal form as a defense evasion technique. When converted back to ASCII, it reveals the command $s/[[:blank:]]\textbackslash{}+//g$, which is used to clean up injected spaces in obfuscated shellcode designed to evade detection.
In Figure 16(c), the string 2E2F676C20262E2F67...6C2026 represents a base64 obfuscated command-line for defense evasion. When decoded the content, it reveals the command $gtcache~\&\textgreater{}/dev/null~\&$, which runs a backdoor process in the background, suppressing output and ensuring persistent session activity without logging.
$\blacksquare$ How well does DefendCLI in the attack correlation with the low-frequency event compared to Nodlink? Figures 16(b), 16(d), 16(e), and 16(f) illustrate the detection results of both detectors in handling complex attack-chain correlations involving low-frequency events.
IV-H Runtime Performance
Table X presents the execution performance of DefendCLI across diverse datasets, highlighting its efficiency through parallelization. For the E3 data sets, DefendCLI runs on a 64,GB RAM, 16-core virtual machine, processing approximately 5,000 records in 6 seconds using multiprocessing. For industrial-scale, real-time detection tasks (A1, A2, A3), the system is deployed on Microsoft Azure DataLake [21], leveraging Kafka-based streaming and ETL pipelines. Under this setup, DefendCLI achieves end-to-end response times of roughly 2 seconds, including network latency, for high-volume threat detection workloads.
TABLE IX: The Scale of Attack-Related Command-Line Detected by DefendCLI and Nodlink
@@ IV-H Runtime Performance
TABLE X: Runtime Performance